OpenStudy (anonymous):

Identify the slope and y-intercept of the inequality 4x + 2y > 8. slope = -2, y-intercept = 4 slope = 2, y-intercept = 4 slope = -2, y-intercept = 8 slope = -1/2 y-intercept = 4

OpenStudy (anonymous):

slope = -2, y-intercept = 4 slope = -2, y-intercept = 8 slope = 2, y-intercept = 4 slope = -1/2 y-intercept = 4

OpenStudy (ash2326):

Since we have to find the slope and y intercept we'll first rewrite the equation as \[4x+2y=8\] do you get this?

OpenStudy (anonymous):

first one :)

OpenStudy (ash2326):

how did you find?

OpenStudy (anonymous):

calculator

OpenStudy (ash2326):

did you use a graphing calculator? how to find y intercept?